A baseball player is running to second base at 6.94 m/s. When he is 5.10 m from the plate he goes into a slide. The coefficient of kinetic friction between the player and the ground is 0.15, and the coefficient of static friction is 0.23. What is his speed when he reaches the plate?




1
Expert's answer
2021-06-07T09:35:05-0400

As the player starts sliding, its initial kinetic energy starts converting to work against friction, so, the final kinetic energy is smaller.


"\frac12 mv^2=\mu_kmgx+\frac12mu^2,\\\space\\
u=\sqrt{v^2-2\mu_kgx},\\
u=\sqrt{6.94^2-2·0.15·9.8·5.1}=5.76\text{ m/s}."